2.0 litres of the ideal gas law (Xenon) is contained with a container at 5.0 atm and 15.0 degrees. How many moles of Xe gas are present.
1
Expert's answer
2020-11-18T13:43:59-0500

PV = nRT T = 15+273= 288K

n= PV/RT = 5×2 / (0.0821 × 288)

n = 0.423 mole